# Rob A. J. Eyles-Ferris,1 Andrew King,1,2,3 Rhaana L. C. Starling,1 Peter G. Jonker,4 Andrew J. Levan,4 Antonio Martin-Carrillo,5 Tanmoy Laskar,6,4 Jillian C. Rastinejad,7,8 Nikhil Sarin,9,10 Nial R. Tanvir,1 Benjamin P. Gompertz,11,12 Nusrin Habeeb,1 Paul T. O’Brien1 and Massimiliano De Pasquale13

### v1

@@ 3.2 A wide angle, low Lorentz factor jet
This model of the early accretion and a jet launched through super-Eddington accretion can easily account for the precursor X-ray emission observed by EP. In this case, we assume the X-rays originate from some internal process in the jet and this emission dominates over any synchrotron from external shocks. We therefore assume a constant radiative and accretion efficiency in the X-ray band, $\epsilon_{X}$, and jet (half-)opening angle, $\theta_{j}$ and the X-ray luminosity should therefore be roughly proportional to the total fallback rate. While we have no knowledge of the observed light curve, our model predicts a stepped or multi-peaked rise with a period approximately equal to the 2825 s identified by Levan et al. (2025a) and remarkably similar to the behaviour exhibited by XRT 000519 (Jonker et al., 2013) and predicted in MacLeod et al. (2014). We note, however, that we have only very weak constraints on the amount of material stripped, and this periodicity could be smoothed out somewhat over the tens of contributing pericentric passages. We therefore do not include it in Figure 1.
It is difficult to independently predict exactly when the peak in the light curve will occur in this model and we therefore assume it occurs with the third Fermi trigger to coincide with the observed X-ray peak. We note that the final disruption could occur earlier, however, assuming the X-rays are primarily powered by accretion, an enhanced accretion rate could still occur coincident with the third trigger. Due to the low mass of the IMBH, and in direct contrast to the prediction from Levan et al. (2025a) which assumed a low redshift event, the accretion rate is likely to stay highly super-Eddington for months [^4] The redshift and therefore luminosity assumed there were much smaller and therefore significantly closer to the Eddington luminosity,. The same jet internal processes should therefore dominate the X-ray light curve over the first days to months while at late times, there may be a sudden drop as the accretion rate drops below Eddington as observed in other relativistic TDEs [^5] It is difficult to say whether this is being seen in the last Chandra observation as the difference with the expected power law decay is low significance.. However, as this is not apparent in the observed light curve, we assume we are still in the jet dominated phase.
@@
Instead of the ‘frozen-in’ approximation, we use a semi-analytical approximation for the fallback rate which partially takes into account the effects of tidal distortion and will be described fully in a future work [^6] To summarise, the tidal distortion results in the specific energy distribution of the star consisting of the core and two ‘wings’. These are approximated with three Gaussians plus an additional constant inside the wings.. We assume the final disruption dominates the long term accretion rate as the earlier stripping episodes are both quicker and will result in significantly less mass being accreted. In Figure 1, we show that our model of a $0.6~M_{\odot}$ WD and a $10^{4}~M_{\odot}$ IMBH can indeed reproduce the long term evolution of GRB 250702B. In particular, the decay slope is naturally consistent with the independently derived eccrentricity of the orbit, $e=0.945$. However, our model does overpredict the observed luminosity from EP. This is not entirely unexpected, however, and suggests that energy is being lost to some other process rather than being radiated in X-rays. In Section 3.3, we discuss the possibility that this missing energy powers inverse Compton upscattering.
@@
We can also use the assumed radiative efficiency of the model to make inferences about the jet properties. In particular, our model assumes a total efficiency of $0.15$. If we make the conservative assumption that $\epsilon_{X}=0.001$, we find that $\theta_{j}\sim 0.12$ rad and therefore the Lorentz factor is $\Gamma\sim 1/\theta_{j}\sim 8.6$. While our assumptions are reasonable, there are some degeneracies in our model that are difficult to account for. However, it is unlikely that the jet properties will significantly differ - in particular, the value we derive for $\theta_{j}$ is probably towards the lower end. Such a jet is wider than typically observed in long GRBs (e.g. Frail et al., 2001; Yi et al., 2017), although could be an outlier to the population, and comparable to that inferred for the relativistic TDE Swift J1644+57 at early times (Metzger et al., 2012; Zauderer et al., 2013) although we note that Beniamini et al. (2023) find a significantly larger $\theta_{j}$ for that TDE. It therefore seems feasible that the inferred jet could form in such a system. While the Lorentz factor is small, it will still be sufficient to produce a radio counterpart through external shock induced synchrotron although we do not examine it in detail here nor make a direct comparison with the extensive radio observations of GRB 250702B already performed (Sfaradi et al., 2025a; Bright et al., 2025; Sfaradi et al., 2025b; Atri et al., 2025; Alexander et al., 2025; Tetarenko et al., 2025; Balasubramanian et al., 2025; Grollimund et al., 2025; Rhodes et al., 2025; Levan et al., 2025a). By analogy with other relativistic TDEs (e.g. Pasham et al., 2015; Hammerstein et al., 2025), the NIR counterpart will trace thermal emission, possibly from an inflated accretion flow or cooling envelope (e.g. Coughlin & Begelman, 2014; Dai et al., 2018; Eyles-Ferris et al., 2022; Metzger, 2022; Sarin & Metzger, 2024). However, this is likely to have a peak bolometric luminosity $\lesssim 10^{45}$ erg s-1 and follow the X-ray in fading rapidly, as shown by Levan et al. (2025a). The non-detection of significant transient light by Gompertz et al. is therefore also consistent with this model.
@@ 3.3 Stream induced shocks and gamma-ray emission
There is, however, another possible emission component that we have not yet examined. Following the disruption or stripping of the WD, the disrupted material will be stretched into a long debris stream. This stream will orbit the IMBH and, depending on the exact conditions, may intersect with itself relatively promptly (e.g. Darbha et al., 2019; Lu & Bonnerot, 2020; Eyles-Ferris et al., 2024) and can influence formation of the accretion flow (Andalman et al., 2022). These collisions are highly energetic and release ‘free’ kinetic energy from material that will still be accreted. They can also enhance the accretion rate as the shocked material is rapidly driven onto the IMBH (Andalman et al., 2022; Eyles-Ferris et al., 2024). If the final disruption does occur earlier than we assume above, this mechanism may explain why the observed X-ray peak would then be delayed relative to the peak in the fallback rate.
@@
In addition to the stream interacting with itself, the earlier tidal stripping of the WD and the resulting debris streams result in additional and, due to our assumed $\beta=1$ curtailing self-intersections, more probable sources of possible interactions. The individual debris streams will behave similarly and in particular, consecutive streams will be almost identical. If, then, enough amount of mass is stripped in a given passage, the debris stream will be long enough that the tip of the debris stream from the subsequent passage will collide with its tail. Such collisions will occur near pericentre and due to the similarity of the debris streams, will be broadly periodic in accordance with the orbital period of the WD. There is also the possibility that remnants from earlier debris streams can be compressed into the same region of space on a somewhat stochastic basis which may boost the mass and therefore energy of these collisions. There is also the possibility of stream-disk collisions that may release energy in a similar fashion to high luminosities (Steinberg & Stone, 2024).
@@
The energy released by these collisions is likely to vary somewhat but can be roughly constrained using the formalism of Darbha et al. (2019). Near the pericentre, the tidal forces exerted by the black hole will compress the streams (e.g. Luminet & Marck, 1985) to a radius $\sim R_{\rm WD}$ and the density at the surface will therefore be comparable to the original WD i.e. $\rho\sim 3M_{\star}/4\pi R_{\rm WD}^{3}$. The kinetic energy of the shocked material is then given by $\rho v_{s}^{2}$ where $v_{s}$ is the shock velocity which is comparable to the pericentric velocity $v_{s}\sim v_{p}=\left(GM_{\bullet}/r_{p}\right)^{1/2}$. We assume that $r_{p}=r_{t}$ for the $M_{\rm WD}=0.6~M_{\odot}$, $M_{\bullet}=10^{4}~M_{\odot}$ system described above and therefore $v_{p}\sim 0.26c$. The energy is deposited into an optically thin surface layer with thickness $\Delta R\lesssim(cR_{\rm WD}/\rho\kappa v_{p})^{1/2}$, where $\kappa$ is the opacity, and therefore the total energy released thermally is given by
@@
where $l_{c}\sim R_{\rm WD}$. Assuming a standard opacity of $\kappa=0.4$ cm2 g-1, we therefore estimate a luminosity of $\sim 6\times 10^{47}$ erg s-1 emitted over a time 0.03 s derived from the light crossing time of the $R_{\rm WD}$. Assuming a perfect blackbody, the spectrum will peak in the hard X-rays at $\sim 50$ keV.
However, whether two consecutive streams collide will also be affected by precession of the debris streams. This precession naturally arises from the fact that the WD’s orbit is almost certainly initially unaligned with the spin axis of the IMBH and is unlikely to have become aligned despite the substantial number of orbits since it had become bound (e.g. Stone & Loeb, 2012; Franchini et al., 2016; Bonnerot et al., 2021). This is consistent with the spacing of the three gamma-ray flares - substantial stream interactions occur coincident with the gamma-ray flares but only a small number or only weak collisions may occur between the second two (‘GRB 250702B’ and ‘GRB 250702E’). Alternatively, it is possible that less mass was stripped during the relevant passages and therefore the debris streams were not long enough to intersect. This could be due to the effects of tidal distortion but requires numerical simulations that are beyond the scope of this work to confirm.
@@
These collisions may not make a significant contribution to the observed X-ray light curve and despite the relatively high peak energy, are still significantly below the spectrum observed by Fermi. They are too faint by over an order of magnitude to trigger Swift’s Burst Alert Telescope assuming a 15 - 150 keV sensitivity of $3\times 10^{-9}$ erg cm-2 s-1 (e.g. Lien et al., 2016). However, in addition to being observed directly, the emission from the collisions provides a significant field of seed photons that could be upscattered by the jet [^7] The accretion disk and any disk winds may also provide additional sources of seed photons. However, the luminosities of these components and the energies of the resulting photons are unlikely to make a significant contribution. The jet itself may also provide seed photons for synchrotron self-Compton but due to the low peak energy and small $\Gamma$ this may only be a small contribution at gamma-ray energies. We therefore do not examine them further here.. If we make the standard approximation of $\nu\sim\Gamma^{2}\nu_{0}$ (e.g. Rybicki & Lightman, 1979; Longair, 2011), the peak in the inverse Compton spectrum should lie at $\sim 3.9$ MeV. This is broadly consistent with the findings of Oganesyan et al. (2025) although they found the peak to lie slightly higher at 10 MeV or so. This could be due to us underestimating $\Gamma$ by a small factor ($\lesssim 2$).
The nature of the seed photon field produced by these shocks can also help explain the highly variable nature of the gamma-rays. As mentioned above, each thermal flash from these collisions will last $\sim 0.03$ s and the chaotic nature of the debris streams will therefore cause significant short timescale variability. It is also worth noting that the emission region of these flashes is at $\sim r_{t}$. From Figure 2, for the $M_{\rm WD}=0.6~M_{\odot}$, $M_{\bullet}=10^{4}~M_{\odot}$ system that we assume, this is roughly a lightsecond away from the jet. Variability on timescales of a second in the rest frame is therefore entirely reasonable to expect in the gamma-ray light curve.
@@
A final consideration is the luminosity of the gamma-rays. From Neights et al. (2025b), the luminosity of each trigger is a few $10^{51}$ erg s-1, significantly higher than the thermal luminosity produced from the stream-stream collisions. The majority of the energy must therefore come from the relativistic electrons within the jet. While this appears to be unachievable from Figure 1, the kinetic luminosity of the jet, i.e. the energy available for upscattering the photons, is significantly higher than the X-ray luminosity. From the results of Wu et al. (2018), the kinetic efficiency relative to the fallback rate could be as high as 0.05, a factor of 50 greater than derived for the X-rays. The energy budget is therefore more than sufficient to boost the shock photons and produce the gamma-rays. We note also that this loss of energy to this process is entirely consistent with our model’s overprediction of the EP-WXT observed luminosity. In fact, even if a smaller kinetic efficiency of only 0.005 is assumed, the converted excess X-ray energy ($\sim 3.4\times 10^{53}$ erg) is consistent with the total energy emitted as gamma-rays across all three triggers ($\sim 3.7\times 10^{53}$ erg). This stream collision-upscattering origin therefore fully accounts for the gamma-rays and explains features iv and v.

### v2

@@ 3.2 A wide angle, low Lorentz factor jet
This model of the early accretion and a jet launched through super-Eddington accretion can easily account for the precursor X-ray emission observed by EP. In this case, we assume the X-rays originate from some internal process in the jet and this emission dominates over any synchrotron from external shocks. We therefore assume a constant radiative and accretion efficiency in the X-ray band, $\epsilon_{X}$, and jet (half-)opening angle, $\theta_{j}$. The X-ray luminosity should therefore be proportional to the total fallback rate. Our model predicts a stepped or multi-peaked rise with a period approximately equal to the 1390 s (converted to rest frame) identified by Levan25 and remarkably similar to the behaviour exhibited by XRT 000519 (Jonker13) and predicted in Macleod14. We note that this is not readily apparent in the precursor light curve as presented by Li25 but is not necessarily ruled out and further investigation of those data may be indicative of our predicted behaviour. We also note, however, that we have only very weak constraints on the amount of material stripped, and this periodicity could be smoothed out somewhat over the tens of contributing pericentric passages. We therefore do not include the precursor emission our fitting below. Due to the low mass of the IMBH, and in direct contrast to the prediction from Levan25 which assumed a low redshift event, the accretion rate is likely to stay highly super-Eddington for months [^6] The redshift and therefore luminosity assumed in Levan25 were much smaller and therefore significantly closer to the Eddington luminosity,. The same jet internal processes should therefore dominate the X-ray light curve over the first days to months while at late times, there may be a sudden drop as the accretion rate drops below Eddington as observed in other relativistic TDEs [^7] It is difficult to say whether this is being seen in the last Chandra observation as the difference with the expected power law decay is low significance.. However, as this is not apparent in the observed light curve, we assume we are in the jet dominated phase throughout.
Instead of the ‘frozen-in’ approximation, we use a semi-analytical approximation for the fallback rate which partially takes into account the effects of tidal distortion. To briefly summarise, the frozen-in approximation assumes the specific orbital energy distribution of the debris stream is fixed from the time of disruption. However, numerical simulations have shown this not to be the case due to the tidal distortion of the star at the time of its disruption and the self-gravity of the stream. We therefore use a semi-analytical approximation for the specific energy distribution given by
$dm/d\varepsilon=\begin{cases}A_{1}e^{-\frac{(\varepsilon-\varepsilon_{0,1})^{2}}{2\sigma_{1}^{2}}},&{\rm for}~\varepsilon<\varepsilon_{c,1},\\ k_{c}+A_{c}e^{-\frac{(\varepsilon-\varepsilon_{0,c})^{2}}{2\sigma_{c}^{2}}},&{\rm for}~\varepsilon_{c,1}\leq\varepsilon\leq\varepsilon_{c,2},\\ A_{2}e^{-\frac{(\varepsilon-\varepsilon_{0,2})^{2}}{2\sigma_{2}^{2}}},&{\rm for}~\varepsilon>\varepsilon_{c,2},\end{cases}$ (1)
i.e. three Gaussians with an additional constant for the central region. These represent the ‘wings’ and core of the star. We fit this model to the numerical results of (Cufari22a) using a least squares method and present our results in Figure 2. Figure 2 shows that our semi-analytical approximation accurately recreates the structure of the numerical result and is a significant improvement on the frozen-in approximation. While the masses at the extremes of the orbital specific energy are underestimated, the effect of this is negligible over the timescales we focus on. The more circular WD orbit required here will shift this distribution towards the negative with an offset of $(1-e)/(1-e_{\rm crit}^{-})$ where $e_{\rm crit}^{-}=1-\frac{2r_{t}}{r_{p}}\left(\frac{M_{\bullet}}{M_{\rm WD}}\right)^{-1/3}$ is the critical eccentricity (Hayasaki18). At a given time, $t$, material with a specific orbital energy will return to the IMBH,
$\varepsilon(t)=-\frac{R_{\rm WD}}{2}\left(\frac{M_{\bullet}}{M_{\rm WD}}\right)^{2/3}\left(\frac{\sqrt{GM_{\bullet}}}{2\pi}\right)^{-2/3}t^{-2/3}$ (2)
and we can then simply combine Equations 1 and 2 to derive the fallback rate, $\dot{m}$ at time $t$. As we assume the X-rays are powered by internal processes with constant efficiency, we convert the fallback rate to a luminosity as $L_{X}(t)=\epsilon_{T}\dot{m}(t)$ where $\epsilon_{T}=2\epsilon_{X}/(1-\cos\theta_{j})$.
Figure 2: Top: the specific orbital energy distribution of a disrupted WD derived using the frozen-in approximation (black), numerical simulations (Cufari22a, red, C22) and our semi-analytical approximation (blue). An eccentricity of $e=1$ is assumed here. Bottom: the ratio between our semi-analytical approximation and the numerical result of (Cufari22a).
@@
We assume the final disruption dominates the long term accretion rate as the earlier stripping episodes are both quicker and will result in significantly less mass being accreted. We fit our model to the long term X-ray decay using the MCMC sampler emcee v3.1.6 (ForemanMackey13) over 5000 iterations with the first 500 discarded as burn-in. In Figure 1, we show 100 traces from the resulting MCMC chain and demonstrate that our model can indeed reproduce the long term evolution of GRB 250702B. We summarise the posterior distributions in Table 1 and also find that our model naturally converges to peak roughly coincident with the third Fermi trigger. However, our model does overpredict the observed luminosity from EP by a factor of two to ten. This is not entirely unexpected, however, and suggests that energy is being lost to some other process rather than being radiated in X-rays. In Section 3.3, we discuss the possibility that this missing energy powers inverse Compton upscattering.
Table 1: The posterior distributions from our MCMC fit to GRB 250702B’s X-ray light curve.
Parameter Posterior
$\log\left(M_{\bullet}/M_{\odot}\right)$ $4.02\pm 0.63$
$M_{\rm WD}$ $0.45\pm 0.20$
$e$ $0.923\pm 0.027$
$\epsilon_{T}$ $0.50\pm 0.25$
@@
We can also use the assumed radiative efficiency of the model to make inferences about the jet properties. We make the conservative assumption that $\epsilon_{X}=0.001$ (Crumley16 find 0.010.10.01-0.1 while including beaming efficiency) and find that $\theta_{j}=0.09^{+0.04}_{-0.02}$ rad. The Lorentz factor is therefore $\Gamma\sim 1/\theta_{j}=11.2^{+2.5}_{-3.3}$. We note that these values are sensitive to our assumed $\epsilon_{X}$ although our assumption is reasonable. Such a jet is wider than typically observed in long GRBs (e.g. Frail01; Yi17), but could be an outlier to the population, and comparable to that inferred for the relativistic TDE Swift J1644+57 at early times (Metzger12; Zauderer13) although we note that Beniamini23 find a significantly larger $\theta_{j}$ for that TDE. It therefore seems feasible that the inferred jet could form in a WD-IMBH TDE. While the Lorentz factor is small, it will still be sufficient to produce a radio counterpart through external shock induced synchrotron although we do not examine it in detail here nor make a direct comparison with the extensive radio observations of GRB 250702B already performed (GCN40979; GCN40985; GCN41053; GCN41054; GCN41059; GCN41061; GCN41145; GCN41147; GCN41215; Levan25). By analogy with other relativistic TDEs (e.g. Pasham15; Hammerstein25), the NIR counterpart will trace thermal emission, possibly from an inflated accretion flow or cooling envelope (e.g. Coughlin14; Dai18; EylesFerris22; Metzger22; Sarin24). However, this is likely to have a peak bolometric luminosity $\lesssim 10^{45}$ erg s-1 and follow the X-ray in fading rapidly, as shown by Levan25. The non-detection of significant transient light by (Gompertz25) is therefore also consistent with this model.
@@ 3.3 Stream induced shocks and gamma-ray emission
There is, however, another possible emission component that we have not yet examined. Following the disruption or stripping of the WD, the disrupted material will be stretched into a long debris stream. This stream will orbit the IMBH and, depending on the exact conditions, may intersect with itself relatively promptly (e.g. Darbha19; Lu20; EylesFerris24) and can influence formation of the accretion flow (Andalman22). These collisions are highly energetic and release ‘free’ kinetic energy from material that will still be accreted. They can also enhance the accretion rate as the shocked material is rapidly driven onto the IMBH (Andalman22; EylesFerris24).
@@
In addition to the stream interacting with itself, the earlier tidal stripping of the WD and the resulting debris streams result in additional and, due to our assumed $\beta=1$ curtailing self-intersections, more probable sources of possible interactions. The individual debris streams will behave similarly and in particular, consecutive streams will be almost identical. If, then, enough amount of mass is stripped in a given passage, the debris stream will be long enough that the tip of the debris stream from the subsequent passage will collide with its tail. Such collisions will occur near pericentre and due to the similarity of the debris streams, will be broadly periodic in accordance with the orbital period of the WD. However, whether two consecutive streams collide will also be affected by precession of the debris streams which naturally arises from the fact that the WD’s orbit is almost certainly initially unaligned with the spin axis of the IMBH and is unlikely to have become aligned despite the substantial number of orbits since it had become bound (e.g. Stone12; Franchini16; Bonnerot21) [^8] This misalignment may also lead to jet precession which can further contribute to the periodicity (e.g. Teboul23; An25).. This is consistent with the spacing of the three gamma-ray flares - substantial stream interactions occur coincident with the gamma-ray flares but only a small number or only weak collisions may occur between the second two. Alternatively, it is possible that less mass was stripped during the relevant passages and therefore the debris streams were not long enough to intersect. This could be due to the effects of tidal distortion but requires numerical simulations that are beyond the scope of this work to confirm.
@@
The energy released by these collisions is likely to vary somewhat but can be roughly constrained using the formalism of Darbha19999While Darbha19 concentrate on ultra-deep encounters where a single debris stream promptly intersects with itself, the same physics are applicable to similar debris streams interacting with each other.. Near the pericentre, the tidal forces exerted by the black hole will compress the streams (e.g. Luminet85) to a radius $\sim R_{\rm WD}$ and the surface density will therefore be comparable to the original WD i.e. $\rho\sim 3M_{\rm WD}/4\pi R_{\rm WD}^{3}$. The kinetic energy of the shocked material is then given by $\rho v_{s}^{2}$ where $v_{s}$ is the shock velocity which is comparable to the pericentric velocity $v_{s}\sim v_{p}=\left(GM_{\bullet}/r_{p}\right)^{1/2}$. We assume that $r_{p}=r_{t}$ and therefore for our derived posteriors, $v_{p}=0.23^{+0.21}_{-0.12}~c$. The energy is deposited into an optically thin surface layer with thickness $\Delta R\lesssim(cR_{\rm WD}/\rho\kappa v_{p})^{1/2}$, where $\kappa$ is the opacity, and therefore the total energy released thermally is given by
@@
where $l_{c}\sim R_{\rm WD}$. Assuming a standard opacity of $\kappa=0.4$ cm2 g-1, we therefore estimate a luminosity of $4.3^{+9.5}_{-3.2}\times 10^{47}$ erg s-1 emitted over a time $\sim 0.03$ s derived from the light crossing time of the $R_{\rm WD}$. Assuming a perfect blackbody, the spectrum will peak in the medium to hard X-rays at $15.9^{+7.7}_{-6.0}$ keV.
@@
These collisions may not make a significant contribution to the observed X-ray light curve and despite the relatively high peak energy, are still significantly below the spectrum observed by Fermi. They are too faint by over an order of magnitude to trigger Swift’s Burst Alert Telescope assuming a 15 - 150 keV sensitivity of $3\times 10^{-9}$ erg cm-2 s-1 (e.g. Lien16). However, in addition to being observed directly, the emission from the collisions provides a significant field of seed photons that could be upscattered by the jet [^10] The accretion disk and any disk winds may also provide additional sources of seed photons. However, the luminosities of these components and the energies of the resulting photons are unlikely to make a significant contribution. The jet itself may also provide seed photons for synchrotron self-Compton but due to the low peak energy and small $\Gamma$ this may only be a small contribution at gamma-ray energies. We therefore do not examine them further here.. If we make the standard approximation of $\nu\sim\Gamma^{2}\nu_{0}$ (e.g. Rybicki_and_Lightman; Longair), the peak in the inverse Compton spectrum should lie at $2.0^{+2.4}_{-1.4}$ MeV, broadly consistent with the $E_{p}$ identified in the spectral fits of Neights25. This could be due to us underestimating $\Gamma$ by a small factor ($\lesssim 2$). The nature of the seed photon field produced by these shocks can also help explain the highly variable nature of the gamma-rays. As mentioned above, each thermal flash from these collisions will last $\sim 0.03$ s and the chaotic nature of the debris streams will therefore cause significant short timescale variability. It is also worth noting that the emission region of these flashes is at $\sim r_{t}$. For a $M_{\rm WD}=0.45~M_{\odot}$, $M_{\bullet}=10^{4}~M_{\odot}$ system that we assume, this is roughly a lightsecond away from the jet. Variability on timescales of a second in the rest frame is therefore entirely reasonable to expect in the gamma-ray light curve.
@@
A final consideration is the luminosity of the gamma-rays. From Neights25, the luminosity of each trigger is a few $10^{51}$ erg s-1, significantly higher than the thermal luminosity produced from the stream-stream collisions. The majority of the energy must therefore come from the relativistic electrons within the jet. While this appears to be unachievable from Figure 1, the kinetic luminosity of the jet, i.e. the energy available for upscattering the photons, is significantly higher than the X-ray luminosity. From the results of Wu18, the kinetic efficiency relative to the fallback rate could be as high as 0.05, a factor of 50 greater than derived for the X-rays. The energy budget is therefore more than sufficient to boost the shock photons and produce the gamma-rays. We note also that this loss of energy to this process is entirely consistent with our model’s overprediction of the EP-WXT observed luminosity. In fact, even if a smaller kinetic efficiency of only 0.005 is assumed, the converted excess X-ray energy ($\gtrsim 10^{54}$ erg) is comparable to the total energy emitted as gamma-rays across all three Fermi-GBM triggers ($\sim 2\times 10^{54}$ erg). This stream collision-upscattering origin therefore fully accounts for the gamma-rays and explains features iv and v.